OpenStudy (anonymous):

MEDAL WILL BE GAVE + FAN!!!! Which fraction is equivalent to 5/24 A. 10/29 B. 15/ 72 C. 25/48 D. 21/100

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Yeah it is B.

OpenStudy (anonymous):

B

OpenStudy (anonymous):

stOOOOOp giving out the answer, EXPLAIN !!!!!!

OpenStudy (anonymous):

metal?

OpenStudy (anonymous):

5/24=.20833333333ect. 10/29= 34482ect. 15/72= 2083333ect. 25/48= 52083333ect. 21/100= .21 The answer is B because it matches the question.

OpenStudy (anonymous):

It's "Medal" if you mean it like a badge. Metal, is something else, it's Iron, Magnesium Copper, Gold and etc. .